### v1

# Adaptive Coordinators and Prompts on Heterogeneous Graphs for Cross-Domain Recommendations

Abstract.
In the online digital world, users frequently engage with diverse items across multiple domains (e.g., e-commerce platforms, streaming services, and social media networks), forming complex heterogeneous interaction graphs. Leveraging this multi-domain information can undoubtedly enhance the performance of recommendation systems by providing more comprehensive user insights and alleviating data sparsity in individual domains. However, integrating multi-domain knowledge for the cross-domain recommendation is very hard due to inherent disparities in user behavior and item characteristics and the risk of negative transfer, where irrelevant or conflicting information from the source domains adversely impacts the target domain’s performance. To address these challenges, we offer HAGO, a novel framework with Heterogeneous Adaptive Graph coOrdinators, which dynamically integrate multi-domain graphs into a cohesive structure by adaptively adjusting the connections between coordinators and multi-domain graph nodes, thereby enhancing beneficial inter-domain interactions while mitigating negative transfer effects. Additionally, we develop a universal multi-domain graph pre-training strategy alongside HAGO to collaboratively learn high-quality node representations across domains. To effectively transfer the learned multi-domain knowledge to the target domain, we design an effective graph prompting method, which incorporates pre-trained embeddings with learnable prompts for the recommendation task. Our framework is compatible with various graph-based models and pre-training techniques, demonstrating broad applicability and effectiveness. Further experimental results show that our solutions outperform state-of-the-art methods in multi-domain recommendation scenarios and highlight their potential for real-world applications. We will open our source code [^1] This paper is currently under review in a conference, we have submitted the source code to the venue. The source link contains the venue information so we will release it after acceptance..
Cross-Domain Recommendation, Graph Prompting, Graph Neural Networks
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Figure 1. The schematic diagram of the existing cross-domain recommendation methods and our perspective.
In real-world applications, users often interact with various items and other users (Li et al., 2024b) across multiple domains within the platform, such as different product categories in e-commerce or distinct content types in streaming services, forming complex heterogeneous interaction graphs. However, the distribution of data across these domains is often uneven. Some domains benefit from abundant user interaction data, while others suffer from data sparsity. How to effectively utilize the abundant interaction data from well-established source domains to enhance recommendation performance in the target domain that faces data sparsity or holds greater strategic importance to the platform (a.k.a cross-domain recommendation) has become an important area in academia and enterprise.
In recent years, various model-centric methods have been proposed to handle this problem yet still leave a huge gap for further improvement. As illustrated in Figure 1(a), one kind of solution is to build some shared-bottom architectures by utilizing shared representations (Singh and Gordon, 2008; Hu et al., 2018) or representation fusion (Zhu et al., 2019; Gao et al., 2019) to represent and transfer information across different domains. The other direction, as shown in Figure 1(b), is to use some mapping functions or transfer layers to map user or item representations from the source domains to the target domain for information enhancement (Man et al., 2017; Zhao et al., 2020; Kang et al., 2019; Guo et al., 2023). Aligned with these two kinds of model-centric frameworks, they also introduce some advanced training skills such as contrastive learning, to promote cross-domain transfer learning further (Li et al., 2024a; Hao et al., 2024; Xie et al., 2022; Liu et al., 2024). Although these methods have achieved notable success, some later theoretical studies (Cao et al., 2023) have found that the performance ceiling of model-centric methods is strictly limited by the intrinsic discrepancy among various domains, indicating that we might not further improve these methods unless we can find some data-level approaches to narrow down the natural gap among various graph domains (as presented in Figure 1(c)).
However, achieving this goal is never easy because we need to overcome at least three tough challenges: Challenge 1, designing a hand-crafted strategy to align different graph domains is far-fetched and tough. Interaction graph data from different domains may exhibit entirely distinct structural patterns and various types of item nodes, resulting in diverse topologies. This makes manually designed strategies far from effective and general. To this end, we need some trainable approaches to learn the potential latent data operation strategy to align these domains yet preserve their important knowledge. Challenge 2, how to integrate and learn the rich knowledge from multiple source domains is tough. The complexity and diversity of these graphs not only enrich the underlying knowledge but also significantly increase the difficulty for models to capture common patterns across multiple domains. Consequently, effectively learning useful knowledge from multi-domain source graphs presents an intricate challenge, as it is crucial for enabling models to capture cross-domain correlations among entities better. Challenge 3, how to transfer the preserved knowledge to the target is still an open problem, and currently often undergoes serious negative transfer issues. Source domains encompass a vast array of information, much of which may not be pertinent or valuable to the target domain. Indiscriminately transferring knowledge from source domains to the target domain may adversely affect the model’s performance, potentially resulting in a decline in recommendation accuracy.
Inspired by the recent success of graph prompting in its powerful data operation capability (Zhao et al., 2024; Sun et al., 2023a, c), we go beyond the previous model-centric paradigm and hope to bridge gaps between diverse graph datasets in a data-centric perspective. We offer HAGO framework, a data-level framework with Heterogeneous Adaptive Graph coOrdinators to address the aforementioned challenges. Specifically, to address Challenge 1, we revisit the theoretical basis of graph prompts in data manipulations, and then we propose graph coordinators to integrate multi-domain graphs in the pre-training stage, which shares a similar theoretical basis as graph prompts. In this way, we can get rid of hand-crafted strategies and learn a potential way to unify various source domains into a unified structure, enabling collaborative multi-domain pre-training to facilitate representation alignment among them. Notably, our graph coordinators are designed to be heterogeneous, using user/item coordinators to connect user/item nodes in graphs, thus preventing negative transfer caused by information confusion between different types of nodes. To address Challenge 2, we give the details of a multi-domain pre-training framework with our graph coordinator mechanism. This framework is model-agnostic and can be easily used in various pre-training strategies and backbones. To address Challenge 3, the coordinators in HAGO connect the nodes in corresponding graphs by edges with adaptive weights, dynamically regulating knowledge transfer both within and across domains. Moreover, we design an efficient graph prompting method that integrates the pre-trained embeddings and prompts for recommendation tasks. We conduct extensive experiments on two real-world platforms with seven different domains, results from which further demonstrate the effectiveness of our method. In summary, the main contributions of our work can be summarized as follows:
  • We introduce HAGO, an effective multi-domain alignment approach with heterogeneous adaptive graph coordinators to dynamically integrate multi-domain graphs into a unified structure. HAGO enhances interconnections between multiple domains while mitigating negative transfer by adaptively adjusting the edge weights between coordinators and nodes.
  • We develop a universal pre-training framework that can integrate various self-supervised learning algorithms alongside HAGO to collaboratively learn high-quality node representations across multiple domains.
  • We design a graph prompting approach to effectively transfer the pre-trained multi-domain knowledge to the target domain. By incorporating pre-trained embeddings with prompts, we improve recommendation accuracy in the target domain.
  • We conduct extensive experiments using seven domains on two platforms. Experimental results demonstrate the superior effectiveness and wide compatibility of HAGO.
2. Background
In this section, we introduce our motivation by comparing two perspectives of cross-domain recommendation. Then, the formal objective of this paper will be given. We give a more comprehensive review of existing related work in Appendix A.1.
Cross-domain Recommendation: A Model-centric Perspective. A mainstream approach for cross-domain recommendation is to employ a shared-bottom architectures (Singh and Gordon, 2008; Hu et al., 2018; Zhu et al., 2019; Gao et al., 2019; Zhu et al., 2024) and mapping functions to project knowledge from source domains to the target (Man et al., 2017; Zhao et al., 2020; Kang et al., 2019; Guo et al., 2023). These approaches are inherently model-centric. Although much progress achieved, some researchers recently have found that model-centric approaches have an unbreakable ceiling caused by the intrinsic gap of the cross-domain data (Cao et al., 2023). That means no matter how better the models are, their upper bound performance is always limited by the data natural gap. Thus, it is imperative to return to the intrinsic nature of the problem and consider how to further break up the ceiling. Intuitively, if we can find some way to narrow the data gap among various domains, we will have a higher performance improvement margin for existing models. However, traditional data dealing approaches heavily rely on manpower and it is far from effective and efficient to manually design data operation strategies for this vision.
Figure 2. Illustration of the proposed HAGO, using two source domains and a target domain as examples. There is partial overlap among user nodes across different domains. For the sake of clarity in the diagram, we omit the overlapping user nodes that are present in all domains. HAGO employs heterogeneous adaptive graph coordinators to organize the multi-domain pre-training collaboratively, and then leverages graph prompts to transfer the knowledge gained from pre-training to the target domain, thereby enhancing downstream recommendation tasks.
Our Motivation: An Attempt of Data-centric Perspective. With the above discussion, this paper tries to explore an enhanced direction, which aims to achieve representation alignment and effective knowledge transfer from a data-centric perspective. Unlike other linear-structured data such as images and text, graph data is highly complex and possesses diverse non-linear structural properties. The nodes in a graph are not isolated; they are deeply influenced by their local neighbors. The connectivity among nodes serves as the carrier of information flow within the graph, reflecting the core characteristics of graph data. Consequently, capturing the common patterns of complex and diverse local structures across multiple graphs to align graph data at the data level particularly presents a more challenging problem. It is infeasible to leverage manually designed strategies to heal the dataset gap effectively. Luckily, recent progress on graph prompting has demonstrated the potential effectiveness in manipulating diversified datasets to be unified (Wang et al., 2024). Graph prompting has the potential to bridge the gaps between multi-domain pre-training and knowledge transfer. With this in mind, we aspire to harness the power of graph prompting to amalgamate graphs from various domains into a cohesive structure, fostering alignment and mitigating negative transfer at the data level.
Formalized Objective. We treat the given $M$ source domains data as user-item interaction graphs, denoted as $\mathcal{G}^{(k)}=(\mathcal{V}^{(k)},\mathcal{E}^{(k)})$, $k=1,2,\cdots,M$. $\mathcal{V}^{(k)}=\mathcal{U}^{(k)}\cup\mathcal{I}^{(k)}$, i.e., the union of the users set $\mathcal{U}^{(k)}$ and items set $\mathcal{I}^{(k)}$ in the $k$-th source domain, and $\mathcal{E}^{(k)}$ is the edge set corresponding to the adjacency matrix $\boldsymbol{A}^{(k)}\in\mathbb{R}^{\mathcal{V}^{(k)}\times\mathcal{V}^{(k)}}$ which records all the interactions between users and items in the $k$-th source domain. The objective is to sufficiently utilize the rich information from source domains $\{\mathcal{G}^{(1)},\cdots,\mathcal{G}^{(k)},\cdots,\mathcal{G}^{(M)}\}$ to improve the Top-N recommendation performance in the target domain $\mathcal{G}^{(t)}=(\mathcal{V}^{(t)},\mathcal{E}^{(t)})$ with the adjacency matrix $\boldsymbol{A}^{(t)}$.
3. Reconcile Multi-Domain Incompatibility via Graph Coordinators and Prompts
Overview: With the above objective, we can see that we should not only try to narrow the gap between the source domains and the target domain, but also need to heal the incompatibility among multiple source domains. To this end, we proposed HAGO as depicted in Figure 2. Specifically, our framework initially employs heterogeneous adaptive graph coordinators to integrate multi-domain graphs into a unified structure and leverages self-supervised learning algorithms for collaborative multi-domain pre-training. Subsequently, graph prompting is strategically implemented to transfer the multi-domain knowledge acquired from pre-training to the target domain, aiming to improve the accuracy of recommendation.
3.1. A Nutshell of the Backbone Model
Notably, our proposed method, HAGO, is a model-agnostic framework that is compatible with various GNN backbones. Here we briefly give a nutshell of LightGCN (He et al., 2020) to capture graph relations, which is widely adopted in recommendation systems.
Entity Embedding: To capture the complex interaction patterns between users and items, we map the ID of each user or item in the given sets of users $\mathcal{U}$ and items $\mathcal{I}$ to a learnable dense vector. This operation utilizes the one-hot vector corresponding to the user or item ID to index the embedding lookup table, which can be formalized as:
(1) $\boldsymbol{e}_{u}=\boldsymbol{E}^{T}_{u}\boldsymbol{o}_{u},\boldsymbol{e}_{i}=\boldsymbol{E}^{T}_{i}\boldsymbol{o}_{i},$
where $\boldsymbol{E}_{u}\in\mathbb{R}^{|\mathcal{U}|\times d}$ and $\boldsymbol{E}_{i}\in\mathbb{R}^{|\mathcal{I}|\times d}$ denotes the embedding lookup table of users and items, and $\boldsymbol{o}_{u}\in\mathbb{R}^{|\mathcal{U}|},\boldsymbol{o}_{i}\in\mathbb{R}^{|\mathcal{I}|}$ are the corresponding one-hot vector of user/item IDs. $|\mathcal{U}|$ and $|\mathcal{I}|$ indicate the numbers of users and items, and $d$ means the embedding size.
Propagation Layer: The graph aggregation process of LightGCN can be formalized as:
(2) $\boldsymbol{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}\boldsymbol{e}_{i}^{(l)},$
$\boldsymbol{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{u}\right|}}\boldsymbol{e}_{u}^{(l)},$
where $\boldsymbol{e}_{u}^{(l)}$ and $\boldsymbol{e}_{i}^{(l)}$ mean the latent representations of user $u$ and item $i$ after $l$-layer graph convolutions, and $\boldsymbol{e}^{(0)}_{u}=\boldsymbol{e}_{u}$, $\boldsymbol{e}^{(0)}_{i}=\boldsymbol{e}_{i}$. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the set of items that are interacted by user $u$ and the set of users that interact with item $i$, respectively.
3.2. Heterogeneous Adaptive Graph Coordinator
3.2.1. Heterogeneous Graph Coordinators
The interaction data is typically modeled by constructing heterogeneous user-item bipartite graphs, which mainly include two types of nodes: users and items, as well as the user-item edges. Naturally, we introduce a set of user coordinators $\mathcal{C}_{u}^{(k)}=\{c_{u,1}^{(k)},c_{u,2}^{(k)},\cdots,c_{u,n}^{(k)}\}$ and item coordinators $\mathcal{C}_{i}^{(k)}=\{c_{i,1}^{(k)},c_{i,2}^{(k)},\cdots,$ $c_{i,n}^{(k)}\}$ for graph $\mathcal{G}^{(k)}$. The whole set of graph coordinators is $\mathcal{C}=\bigcup_{k}\mathcal{C}_{u}^{(k)}\cup\bigcup_{k}\mathcal{C}_{i}^{(k)}$, and the number of coordinators is $|\mathcal{C}|=2n$. Each graph coordinator connects to the nodes in the corresponding graph through dynamically weighted direct edges, and corresponds to a learnable embedding which can be obtained by retrieving the coordinator embedding table $\boldsymbol{E}_{c}\in\mathbb{R}^{|\mathcal{C}|\times d}$ as the same way in Equation 1.
3.2.2. Adaptive Connections between Coordinators and Graphs
To enhance the effectiveness of cross-domain knowledge transfer while alleviating the negative transfer phenomenon, we propose an adaptive edge weight learning strategy to dynamically regulate the strength of information transfer between the coordinators and the nodes. Our proposed strategy encompasses both soft/hard cases and can be formalized as follows:
(3) $w=\left\{\begin{aligned} \frac{\boldsymbol{e}\cdot\boldsymbol{e}_{c}}{||\boldsymbol{e}||||\boldsymbol{e}_{c}||},&\ \ \textit{if}\ \ \boldsymbol{e}\cdot\boldsymbol{e}_{c}>0,\qquad\textit{(Soft Case)}\\ 0\qquad,&\ \ \textit{if}\ \ \boldsymbol{e}\cdot\boldsymbol{e}_{c}\leq 0,\qquad\textit{(Hard Case)}\end{aligned}\right.$
where $\boldsymbol{e}_{c}$ and $\boldsymbol{e}$ are the embeddings of the coordinator and node, and $w$ is the edge weight between them. ()($\cdot$)( ⋅ ) means the inner product operation between two vectors. In the soft case, the edge weights are determined by the cosine similarity between the embeddings of the coordinator and the node; in the hard case, there is no connection between the corresponding coordinator and node. Consequently, the edge weights range between 0 and 1. Following this strategy, the model can capture the degree of association between the coordinators and the nodes through embedding learning, thereby adaptively determining whether there is an edge connection and adjusting the strength of the connection.
3.2.3. Interconnections between Coordinators
Through the aforementioned strategy, we establish adaptive linkages between graph coordinators and their respective graphs, empowering these coordinators to bolster individual graph models by regulating the flow of information among the nodes. Our vision, however, transcends this capability, as we aspire to model collaborative connections across multiple domains within our dataset to facilitate multi-domain representation alignment. Consequently, we establish direct cross-domain interconnections between the graph coordinators associated with each domain graph, effectively amalgamating the entire collection of graphs into a cohesive structure. These graph coordinators assume the pivotal role of information exchange hub between graphs, with the connecting edges between coordinators acting as bridges for this communication. Notably, our proposed HAGO framework includes distinct coordinators for users and items, facilitating connections solely between coordinators of different types. This structured relationship can be effectively depicted using an adjacency matrix $\boldsymbol{R}_{C}$:
(4) $\begin{array}[]{cc}\boldsymbol{R}_{C}=\left[\begin{array}[]{cc}\boldsymbol{0}_{|\mathcal{C}_{u}|\times|\mathcal{C}_{u}|}&\boldsymbol{J}_{|\mathcal{C}_{u}|\times|\mathcal{C}_{i}|}\\ \boldsymbol{J}_{|\mathcal{C}_{i}|\times|\mathcal{C}_{u}|}&\boldsymbol{0}_{|\mathcal{C}_{i}|\times|\mathcal{C}_{i}|}\end{array}\right],&\begin{array}[]{cc}\leftarrow user\\ \leftarrow item\end{array}\\ \begin{array}[]{cc}\uparrow&\qquad\uparrow\\ user&\qquad item\end{array}&\end{array}$
where $\boldsymbol{0}_{p\times q}$ and $\boldsymbol{J}_{p\times q}$ are zero matrix and one matrix with shape of $(p,q)$.
3.2.4. Multi-domain Pre-training
The graph coordinators we introduce serve as a unifying thread, weaving together the diverse domains of our dataset into an integrated tapestry through their strategic architectural integration. Constructing upon this foundation, we introduce a universal multi-domain graph pre-training framework to collaboratively refine high-quality node representations while facilitating representation alignment. By integrating node-level pre-training strategies exemplified by GRACE (Zhu et al., 2020), which shows strong effectiveness in learning embeddings within individual graphs through self-supervised learning, with heterogeneous adaptive graph coordinators, we achieve a globally unified multi-domain pre-training framework. GRACE (Zhu et al., 2020) uses graph augmentation techniques such as edge removing and feature masking to generate pairs of positive samples, and then it employs the InfoNCE (van den Oord et al., 2018) loss function to ensure that these positive samples are brought close together while negative samples are pushed apart.
In order to demonstrate the wide versatility of our framework, we subsequently adopt other state-of-the-art graph pre-training methods like BGRL and COSTA, integrating them into our multi-domain pre-training framework, in Section 4.4. COSTA builds upon GRACE by incorporating the novel Covariance-Preserving Feature Augmentation technique, which effectively tackles the bias challenges present in the original augmentation strategies. BGRL introduces the concept of Bootstrapping Latent Loss, which is designed to align the node representations produced by the online and target encoders with different updating strategies.
3.3. Graph Prompting for Transferring
Prompting technology has been widely used in the NLP area, particularly in achieving general task performance of Large Language Models (Shu et al., 2024). Inspired by this, graph prompting offers some prompt nodes within the graph structure or the enhancement of node features with prompt tokens, and has been widely demonstrated as a powerful method to learn various potential data operations  (Fang et al., 2023; Zhao et al., 2024; Sun et al., 2023a; Wang et al., 2024). This makes graph “pre-training and prompting” stand out in transferring knowledge for graphs compared with traditional “pre-training and fine-tuning” (Li et al., 2024c). Next, we introduce our designed graph prompts customized for the cross-domain recommendations.
3.3.1. Graph Prompt
The prompt embeddings of users and items in the target domain are integrated with the pre-trained embeddings of nodes from multi-domain pre-training to perform the recommendation task. The input features of the model after adding graph prompts can be represented as:
(5) $\boldsymbol{X}^{\#}=\{\boldsymbol{e}^{*}_{u_{1}}+\boldsymbol{p}_{u_{1}},\cdots,\boldsymbol{e}^{*}_{u_{|\mathcal{U}|}}+\boldsymbol{p}_{u_{|\mathcal{U}|}},\boldsymbol{e}^{*}_{i_{1}}+\boldsymbol{p}_{i_{1}},\cdots,\boldsymbol{e}^{*}_{i_{|\mathcal{I}|}}+\boldsymbol{p}_{i_{|\mathcal{I}|}}\},$
where $\boldsymbol{e}^{*}_{u},\boldsymbol{e}^{*}_{i}$ are the node embeddings of the user $u$ and item $i$ pre-trained from the multi-domain pre-training organized by HAGO, and $\boldsymbol{p}_{u},\boldsymbol{p}_{i}$ denote the corresponding prompt embedding. These pre-trained embeddings contain a wealth of multi-domain knowledge absorbed during the pre-training phase, which can effectively enhance the model’s recommendation accuracy in the target domain. Graph prompts serve as the bridge to align the pre-trained node embeddings to the downstream task in the target domain.
3.3.2. Training Objective
By utilizing the graph convolutional network as outlined in Section 3.1 to operate on the unified multi-domain graph framework by the graph coordinators, we derive the final node representations as follows:
(6) $\boldsymbol{o}_{u}=\sum\limits_{l=0}^{L}\alpha_{l}\boldsymbol{e}_{u}^{(l)},\boldsymbol{o}_{i}=\sum\limits_{l=0}^{L}\alpha_{l}\boldsymbol{e}_{i}^{(l)},$
where $\boldsymbol{o}_{u},\boldsymbol{o}_{i}$ denote the final representations of user $u$ and item $i$. $\alpha_{l}$ is the weight of the representation after $l$-layers which satisfies $\sum\limits_{l=0}^{L}\alpha_{l}=1$, where $L$ is the number of layers in the model.
We leverage the Bayesian Personalized Ranking (BPR) loss function (Rendle et al., 2009), a prevalent choice for pairwise ranking within the recommendation research area, to train our model’s performance on the target domain’s recommendation tasks. The BPR loss leads the model to assign higher predictive scores to user-item interactions that have occurred compared to those that have not. Thus, the final training objective can be articulated as follows:
(7) $\mathcal{L}$ $=\mathcal{L}_{BPR}+\mathcal{L}_{Reg}$
$=-\frac{1}{|\mathcal{B}|}\sum\limits_{(u,i,j)\in\mathcal{B}}\mathrm{ln}\sigma(\hat{y}_{ui}-\hat{y}_{uj})+\lambda||\Theta||^{2},$
where the second term is the $2$-regularization for potential overfitting issues. $(u,i,j)$ from batch $\mathcal{B}$ is a triple including user $u$ and item $i,j$. The interaction between user $u$ and item $i$ represents a positive pair, while the absence of interaction between user $u$ and item $j$ signifies a negative pair. $\hat{y}_{ui}=\boldsymbol{o}_{u}\cdot\boldsymbol{o}_{i}$ and $\hat{y}_{uj}=\boldsymbol{o}_{u}\cdot\boldsymbol{o}_{j}$ are the predicted value of interaction occurring between user $u$ and item $i,j$. $\Theta$ denotes the entire ensemble of trainable parameters that the model employs and $\lambda$ controls the regularization degree.
3.4. Why It Works?
Theoretical Support: Our proposed framework shares a common theoretical bedrock with graph prompts designed for knowledge transfer. The primary difference is that while graph prompts are generally targeted at downstream tasks, we utilize the term ”graph coordinators” for the pre-training phase. Studies (Wang et al., 2024) have shown that for a given graph $\mathcal{G}$, we can find some potential graph transformation $T(\cdot)$ to translate the original graph as a new one like $T(\mathcal{G})$ such that the pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ can approximate to an optimal function $C$ for the target domain task, which can be described as:
(8) $f_{\theta^{*}}(T(\mathcal{G}))=C(\mathcal{G}).$
It can be proven that graph prompts can be treated as a learnable method to find such data transformation (Fang et al., 2023; Sun et al., 2023a). Specifically, given a pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ with fixed parameters $\theta^{*}$ and learnable graph prompt $P(\cdot)$, it can be shown that:
(9) $f_{\theta^{*}}(P(\mathcal{G}))=f_{\theta^{*}}(T(\mathcal{G}))+O_{fp},$
where $O_{fp}$ is the bounded error associated with $f_{\theta^{*}}(\cdot)$ and $P(\cdot)$. In other words, graph prompts have the capability to be equivalent to certain graph transformation operations that enable the transformed graph, once processed by the pre-trained graph encoder, to produce optimal embeddings. In this paper, our heterogeneous adaptive graph coordinators can be treated as a variant of graph prompts, which is used in the pre-training stage, and thus share the same foundation for data operations. Based on this idea, we give a data-centric solution to heal the gap among multiple source domains and the target domain such that the performance ceiling of knowledge transfer can be further improved as mentioned in our motivation (Section 2).
Compatibility Analysis: Our proposed HAGO is model-agnostic because it extends the graph structure without regarding to message passing and aggregation within the graph, making it compatible with various graph-based backbone networks. Graph Prompting can also be seamlessly integrated with various graph pre-training algorithms. In our subsequent experiments, comprehensive compatibility analysis demonstrates the broad compatibility and wide effectiveness of HAGO with different backbone networks and pre-training methods.
Complexity Analysis: The time complexity of our proposed HAGO framework is intrinsically linked to that of the backbone model. The additional time complexity introduced by HAGO primarily stems from the heterogeneous adaptive graph coordinators and their connections with the multi-domain graphs. Our backbone model, LightGCN, has a time complexity of $\mathcal{O}(|\mathcal{E}^{(t)}|Ld)$, where $|\mathcal{E}^{(t)}|$ represents the number of edges in the target-domain graph, and $L$ and $d$ denote the number of layers and the dimension of the network, respectively. Consequently, for HAGO, the time complexity during the multi-domain pre-training phase is $\mathcal{O}((\sum\limits_{k=1}^{M}|\mathcal{E}^{(k)}|+|\mathcal{E}^{(t)}|+|\mathcal{E}_{C}|)Ld)$ where $|\mathcal{E}_{C}|$ is the number of edges introduced by HAGO; and when transferring to downstream recommendation tasks via graph prompts, the time complexity remains $\mathcal{O}(|\mathcal{E}^{(t)}|Ld)$.
4. Evaluation Analysis
4.1. Experiment Settings
Table 1. Dataset Statistics
dataset domain #users #items #interactions sparsity
Douban Music 15,996 39,749 1,116,984 99.82%
Movie 22,041 25,802 2,552,305 99.55%
Book 18,086 33,068 809,248 99.86%
Amazon Book 135,110 115,173 4,042,382 99.97%
Electronic 112,988 41,905 1,880,865 99.96%
Movie 26,969 18,564 762,957 99.85%
Toy 11,687 8,289 202,230 99.79%
Table 2. Overall comparison results of our proposed HAGO and state-of-the-art baselines on two real-world multi-domain datasets. The best-performing method is highlighted in bold, while the second-best is indicated with an underline. An asterisk (*) denotes statistical significance (p ¡ 0.05) when comparing HAGO to the best-performing baseline results.
Single-domain Method Cross-domain Method
Dataset Metric NGCF LightGCN XSimGCL CMF EMCDR SSCDR DeepAPF BiTGCF CUT Ours.
Recall@10 0.1195 0.1665 0.1660 0.1143 0.1264 0.1430 0.1442 0.1543 0.1669 0.1689
HR@10 0.2382 0.3116 0.3117 0.2138 0.2498 0.2801 0.2820 0.2879 0.3123 0.3168
NDCG@10 0.0807 0.1210 0.1213 0.0728 0.0870 0.1024 0.1029 0.1103 0.1211 0.1227
Douban MRR 0.0997 0.1516 0.1526 0.0842 0.1079 0.1290 0.1288 0.1362 0.1532 0.1535
Recall@10 0.1064 0.1627 0.1630 0.1195 0.1236 0.1481 0.1542 0.1605 0.1641 0.1662
HR@10 0.1236 0.1902 0.1913 0.1400 0.1399 0.1724 0.1796 0.1874 0.1918 0.1950
NDCG@10 0.0675 0.1141 0.1147 0.0667 0.0834 0.1030 0.1051 0.1131 0.1158 0.1177
Amazon MRR 0.0601 0.1074 0.1088 0.0542 0.0752 0.0951 0.0963 0.1056 0.1055 0.1116
Data Preparation and Setup: We conduct comprehensive experiments on two popular public recommendation datasets: Amazon dataset (Ni et al., 2019) [^2] http://jmcauley.ucsd.edu/data/amazon is gathered from relevant product evaluations on Amazon.com, which are extensively utilized to assess the performance of cross-domain recommender systems. For our study, we have chosen four categories: “Books”, “Electronics”, “Movies and TV”, and “Toys and Games”; Douban dataset (Zhao et al., 2021) [^3] https://recbole.s3-accelerate.amazonaws.com/CrossDomain/Douban.zip is collected from Douban website, which is a large-scale platform for rating and reviewing movies, TV shows, books, and music. This dataset includes three domain datasets: “Douban-Movie”, “Douban-Book”, and “Douban-Music”. In our experiments, we adopt the domain with the least interactions in each dataset as the target domain, namely “Douban-Book” and “Amazon-Toy”, with the other domains serving as the source domains. More detailed statistics of the datasets are given in Table 1.
Evaluation Metrics: We assess the performance of our method by four widely used metrics in recommendations: Recall@10, HR@10 (Hit Ratio), NDCG@10 (Normalized Discounted Cumulative Gain), and MRR (Mean Reciprocal Rank) for top-10 recommendation. Kindly note that we fully rank the entire items within the datasets for calculating the above top-10 recommendation performance rather than evaluate these metrics by pre-sampling a small group of the total items. This is closer to the real industrial applications and raises more challenging tasks for the compared models.
Baselines: We compare our method with NINE state-of-the-art methods, which can be grouped into two categories: (1) Single-domain Recommendation Methods withNGCF (Wang et al., 2019), LightGCN (He et al., 2020), and XSimGCL (Yu et al., 2024). (2) Cross-domain Recommendation Methods with CMF (Singh and Gordon, 2008), EMCDR (Man et al., 2017), DeepAPF (Yan et al., 2019), SSCDR (Kang et al., 2019), BiTGCF (Liu et al., 2020), and CUT (Li et al., 2024a). The details of these methods are present in Appendix A.2.
Implementation Details: All the baselines and our proposed HAGO are implemented based on the PyTorch (Paszke et al., 2019). Specifically, CUT (Li et al., 2024a) is implemented following its official code, and other baselines are developed based on the open-source RecBole library (Zhao et al., 2021). For a fair comparison, we set the embedding size to $64$ for all methods. For all methods that use LightGCN (He et al., 2020) as the backbone, we set the number of layers to $2$ and the weight $\boldsymbol{\alpha}$ of each layer’s representation to $[0,0,1]$, i.e., LightGCN-single in LightGCN paper (He et al., 2020). We conduct a grid search for the key hyperparameters for all baselines. For our proposed HAGO, we perform pretraining using Adam optimizer with a learning rate of $0.005$ and batch size of $1,024$, while a learning rate of $0.1$ and batch size of $4,096$ for training. And the numbers of user and item graph coordinators are both set to $n=5$.
4.2. Overall Performance
The overall performance comparison of our proposed HAGO and state-of-the-art baselines are shown in Table 2. According to the experiment results, we have the following observations:
  • The approaches with LightGCN as the backbone network (i.e., XSimGCL, BiTGCF, CUT, HAGO) demonstrate exceptional performance, which benefits from LightGCN’s strong generalization capabilities and its efficient capturing of intricate, granular collaborative signals. Compared to other backbones, such as Matrix Factorization (MF), LightGCN employs a simple yet effective approach to extract higher-order connectivity from the user-item interaction graph.
  • Cross-domain knowledge transfer does not always result in superior performance. The performance of many cross-domain recommendation methods often falls short when compared to the elite single-domain counterparts. XSimGCL demonstrates superior performance over the majority of its cross-domain peers on both two datasets. This might be caused by that the performance gains derived from cross-domain knowledge transfer pale compared to the substantial advancements achieved by improving the model’s underlying architecture, and the knowledge transfer from the source domain to the target domain in many cross-domain methods may not be effective enough, and it might even result in the negative transfer issue.
  • Our proposed HAGO consistently achieves the best performance across all four metrics on two multi-domain datasets, demonstrating its superior effectiveness. It gains 1.20% and 1.28% improvements on Recall@10, 1.44% and 1.67% on HR@10, 1.32% and 1.64% on NDCG@10, and 0.59% and 2.57% on MRR against the runner-up on Douban and Amazon datasets. HAGO owes its success to employing graph coordinators that seamlessly integrate multi-domain graphs into a cohesive framework for multi-domain collaborative pre-training. These graph coordinators adaptively connect with corresponding domain graphs, dynamically adjusting the information communication within and across domains, achieving information enhancement effectively while mitigating negative transfer. Moreover, it adopts graph prompts to efficiently transfer valuable knowledge from multiple source domains to the target domain, thereby boosting the recommendation performance.
4.3. Effectiveness Analysis on Healing Negative Transfer Problem
To delve into the effect of the graph coordinators in our proposed HAGO framework and to evaluate its capability to mitigate negative transfer, we craft a variety of model variants predicated on distinct coordinator designs:
  • w/o GO: Without Graph Coordinator implies that the graph pre-training process is conducted in an isolated manner across the various domain graphs.
  • HomoGO: This design features homogeneous graph coordinators that establish direct connections, whose edge weight is 1, with all nodes within the associated domain graph, treating user and item nodes uniformly.
  • HeterGO: This design employs heterogeneous graph coordinators that connect exclusively with all user nodes through the user coordinators and with all item nodes through the item coordinators within the respective domain graph.
Table 3. Comparative analysis results of different Graph Coordinator strategies. Orange and cyan indicate performance gains or losses compared to the Backbone, with deeper shades representing greater relative changes.
Douban
Method Recall@10 HR@10 NDCG@10 MRR
Backbone 0.1665 0.3116 0.1210 0.1516
w/o GO 0.1671 0.3121 0.1214 0.1518
HomoGO 0.1667 0.3115 0.1210 0.1515
HeterGO 0.1676 0.3130 0.1217 0.1519
HAGO 0.1689 0.3168 0.1227 0.1535
Amazon
Method Recall@10 HR@10 NDCG@10 MRR
Backbone 0.1627 0.1902 0.1141 0.1074
w/o GO 0.1624 0.1901 0.1139 0.1067
HomoGO 0.1630 0.1901 0.1171 0.1082
HeterGO 0.1646 0.1917 0.1176 0.1110
HAGO 0.1662 0.1950 0.1177 0.1116
The comparative experimental results of different model variants on two datasets are presented in Table 3. For a fair comparison, the number of coordinators is set to be consistent for HomoGO, HeterGO, and HAGO. From the experimental outcomes, we can draw the following observations:
  • We can observe negative transfer phenomena happen in some cases, as indicated by the cyan-shading in the table. In the Amazon dataset, the performance of w/o GO declines compared to the Backbone Model; in the Douban dataset, HomoGO, which employs homogeneous graph coordinators for cross-domain information transfer during the pre-training phase, also experiences a certain degree of performance damage relative to w/o GO. In fact, the potential for HomoGO to lead to negative transfer effects is foreseeable. Homogeneous coordinators in HomoGO that indiscriminately connect to user and item nodes in a heterogeneous interaction graph can lead to information confusion across different types of nodes. Hence, we employ a heterogeneous design for the graph coordinators in HAGO to avoid this conflict.
  • HAGO consistently leads to superior recommendation performance over other model variants across both datasets, securing significant improvements. The success of HAGO can be primarily attributed to two key aspects. Firstly, HAGO integrates multi-domain graphs into a cohesive structure via heterogeneous graph coordinators, effectively avoiding the potential negative transfer caused by information confusion between different types of nodes as seen in HomoGO. This is also substantiated by the consistent improvements in HeterGO results compared to HomoGO. Secondly, HAGO addresses the negative transfer problem effectively by employing an adaptive, trainable connection architecture that links the graph coordinators with multi-domain graphs. This connection structure facilitates dynamic adjustment of the transfer of more valuable information to the target domain.
4.4. Compatibility Analysis with Different Backbones and Pre-training Strategies
Table 4. Compatibility Analysis with Different Backbones. Orange and cyan indicate performance gains or losses compared to the corresponding Backbone, with deeper shades representing greater relative changes.
Douban
Method Recall@10 HR@10 NDCG@10 MRR
GCN 0.1198 0.2392 0.0819 0.1023
HAGO(GCN) 0.1250 0.2446 0.0855 0.1056
GAT 0.1178 0.2315 0.0795 0.0977
HAGO(GAT) 0.1248 0.2416 0.0850 0.1043
LightGCN 0.1665 0.3116 0.1210 0.1516
HAGO(LightGCN) 0.1689 0.3168 0.1227 0.1535
Amazon
Method Recall@10 HR@10 NDCG@10 MRR
GCN 0.1183 0.1444 0.0802 0.0682
HAGO(GCN) 0.1274 0.1506 0.0820 0.0742
GAT 0.1316 0.1539 0.0929 0.0878
HAGO(GAT) 0.1350 0.1571 0.0939 0.0874
LightGCN 0.1627 0.1902 0.1141 0.1074
HAGO(LightGCN) 0.1662 0.1950 0.1177 0.1116
Our proposed HAGO is not confined to a specific model but is designed as a flexible architecture that can be effortlessly combined with a variety of backbone networks serving as graph encoders, as well as with diverse graph pre-training algorithms. To demonstrate the adaptability of HAGO, we conduct comparative experiments by integrating it with other notable graph neural networks, including GCN and GAT. Furthermore, we investigate the incorporation of other self-supervised learning strategies for pre-training the graph, such as BGRL and COSTA.
Table 5. Compatibility Analysis with Different Graph Pretraining Methods. Orange and cyan indicate performance gains or losses compared to the Backbone, with deeper shades representing greater relative changes.
Douban
Method Recall@10 HR@10 NDCG@10 MRR
Backbone 0.1665 0.3116 0.1210 0.1516
Backbone+BGRL 0.1679 0.3145 0.1218 0.1517
HAGO(BGRL) 0.1704 0.3166 0.1234 0.1537
Backbone+COSTA 0.1679 0.3141 0.1213 0.1511
HAGO(COSTA) 0.1698 0.3143 0.1231 0.1523
Amazon
Method Recall@10 HR@10 NDCG@10 MRR
Backbone 0.1627 0.1902 0.1141 0.1074
Backbone+BGRL 0.1642 0.1904 0.1164 0.1104
HAGO(BGRL) 0.1676 0.1953 0.1199 0.1144
Backbone+COSTA 0.1640 0.1918 0.1160 0.1104
HAGO(COSTA) 0.1679 0.1955 0.1181 0.1119
The compatibility analysis experiments of HAGO combined with different backbones and graph pre-training algorithms are presented in Table 4 and 5, respectively. From the experimental results, we can see that HAGO consistently enhances the performance of all backbones, and achieves unified modeling that improves the performance of graph pre-training algorithms on isolated graphs. These results demonstrate the broad compatibility and universal applicability of the HAGO framework we proposed.
4.5. Impact Analysis of Coordinator Number
We conduct comparative experiments around HAGO with different numbers of graph coordinators, and the results of HR@10 and NDCG@10 are presented in Figure 3. Notably, HAGO includes both user and item graph coordinators, and the number of each type of coordinator is $n$. In experiments, we simultaneously vary both the numbers of user and item coordinators. Thus, the total number of graph coordinators in HAGO is $2n$. While examining the impact of the number of coordinators, we keep other hyperparameters fixed.
(a) Douban
(b) Amazon
Figure 3. Performances with different numbers of graph coordinators $2n$.
Observing the results, it is evident that the optimal performance of HAGO is reached when the number of graph coordinators is moderate. When there are too few coordinators, they are inadequate to handle the demanding task of transferring knowledge both within and across domains efficiently. As the number of coordinators increases, the marginal benefits to model performance from the additional coordinators diminish, and there may even be a decline in performance, potentially due to excessive knowledge transfer.
4.6. Visualization Analysis
To investigate whether the HAGO framework we propose can effectively align the representation spaces and facilitate cross-domain knowledge transfer among multiple domains, we conduct a visualization analysis of the pre-trained embeddings for item nodes in each domain, following the analytical methods in (Yu et al., 2022). We randomly sample 2000 items from each of the three domains in the Douban dataset (Douban-Music, Douban-Movie, and Douban-Book). Subsequently, we utilize t-SNE (Van der Maaten and Hinton, 2008) to perform dimensionality reduction to map these item embeddings to two-dimensional feature vectors in the $\mathbb{R}^{2}$ and normalize them onto the unit circle. Employing Gaussian kernel estimation (Botev et al., 2010), we plot the feature distribution and corresponding angle distribution of the item samples from each domain as they are mapped onto the unit circle.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(a) w/o GO
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(b) HAGO
Figure 4. Visualization Analysis of Item Embeddings in Douban Dataset. We perform dimensionality reduction to map item embeddings in three domains of the Douban dataset to two-dimensional feature vectors and normalize them onto the unit circle. We show the distribution of normalized features in the unit circle and the distribution of the corresponding angles, for (a) w/o GO (i.e., isolated pre-training without graph coordinators), and (b) HAGO.
Based on this approach, we perform visualizations for the isolated pre-training without graph coordinators (w/o GO), as well as the proposed HAGO, as shown in Figure 4. From the visualization results, it is observable that the item embeddings learned by the model without GO exhibit a significant clustering effect, with the embeddings from the three domains converging in three distinct directions on the unit circle, and their angle distributions are sharper. It indicates that isolated pre-training without graph coordinators fails to capture the connections between nodes from different domains, resulting in a disjointed embedding space across domains. In contrast, for HAGO, it is evident that the feature distributions corresponding to each domain are more uniformly distributed and interweave with one another. This not only suggests that HAGO can capture higher-quality embeddings that are more evenly spread throughout the embedding space, but also that the interweaving distribution of embeddings from different domains indicates HAGO’s capability to capture cross-domain associations among items. This observation underscores the effectiveness of HAGO in enhancing the expressiveness of embeddings by promoting a more cohesive structure within multi-domain graphs, thereby facilitating knowledge transfer across domains. We give more additional case studies in Appendix A.3.
5. Conclusions
Our work delves into the multi-domain representation alignment and the challenging issue of negative transfer in cross-domain recommendation tasks. To tackle these intricate challenges, we introduce the HAGO framework, designed to integrate multi-domain graphs into a cohesive structure. Within HAGO, graph coordinators dynamically connect with corresponding graphs based on adaptive weights, thereby regulating intra-domain and inter-domain information transfer to mitigate the negative transfer problem. We conduct comprehensive experiments to demonstrate the effectiveness of HAGO and vividly discuss its role in representation alignment and the facilitation of cross-domain transfer through visual analysis and case studies.
Appendix A Appendix
The Appendix is organized as follows: We supplement a more comprehensive review of related work in Appendix A.1, then give a more detailed introduction to our compared baselines in Appendix A.2. In Appendix A.3, we extend our experimental discussion with additional case studies. We open our source code in Appendix A.4.
A.1. Related Work
A.1.1. Self-Supervised Graph Learning
Self-supervised graph learning is a machine learning algorithm that trains graph models using supervision signals generated from the graph data itself, without relying on additional labels (Sun et al., 2023b). Specifically, self-supervised graph learning algorithms can be categorized into generative, predictive, and contrastive approaches  (Liu et al., 2023; Wu et al., 2023). Generative methods typically rely on an encoder-decoder structure to capture generalized latent representations through tasks like feature or structure reconstruction, such as MGAE (Wang et al., 2017) and GALA (Park et al., 2019) for feature reconstruction, and GAE (Kipf and Welling, 2016) and GraphMAE (Hou et al., 2022, 2023) for structure reconstruction. Predictive methods use pseudo-labels generated from the data itself as supervision signals. For instance, S2GRL (Peng et al., 2020) and PairwiseDistance (Jin et al., 2020) use the prediction of the shortest path between two nodes, while M3S (Sun et al., 2020) and SHGP (Yang et al., 2022) generate pseudo-labels through clustering. Contrastive methods, which are the most widely studied, learn representations by constraining positive samples to be closer and pushing negative samples apart. These methods typically augment graph data by adding or removing edges, dropping nodes, feature transformations, or dropping features to generate a series of positive samples, then utilize loss functions such as InfoNCE  (Zhu et al., 2020; You et al., 2020; Xia et al., 2022; Zhu et al., 2022; Zhang et al., 2022), Jensen-Shannon Divergence  (Hassani and Ahmadi, 2020; Sun et al., 2021), or Bootstrapping Latent Loss  (Thakoor et al., 2021) to constrain the distance between positive and negative samples.
A.1.2. Cross-domain Recommendation
Cross-domain recommendation refers to the recommendation algorithms that enhance recommendations in one domain by transferring knowledge from other domains. These algorithms can be categorized into two types: those without overlapping users/items and those with overlapping users/ items. In non-overlapping cases, research mainly focuses on transferring or aligning semantic information across domains to facilitate knowledge transfer between the source and target domains  (Li et al., 2009; Liu et al., 2022). For overlapping cases, studies leverage shared users and/or items to transfer knowledge from the source domains to the target domain. Early attempts utilize shared representations (such as CMF (Singh and Gordon, 2008)) or concatenated representations from different domains (such as DTCDR (Zhu et al., 2019)) to make predictions, learning from both the source and target domains. Later researchers explore building cross-domain connections between neural network models to enhance knowledge transfer, such as CoNet (Hu et al., 2018) and BiTGCF (Liu et al., 2020). EMCDR (Man et al., 2017) is the first to propose a mapping function to transfer knowledge at the latent representation level, which is further expanded by models like CATN (Zhao et al., 2020) and SSCDR (Kang et al., 2019). More recent studies, such as CUT (Li et al., 2024a), have begun exploring the utilization of contrastive learning and other self-supervised techniques to transfer useful knowledge from the source to the target domain, improving recommendation accuracy.
A.2. Details of Baselines
In order to thoroughly assess the effectiveness of our proposed HAGO, we conduct a comprehensive comparison with nine state-of-the-art cross-domain baseline models:
  • Single-domain Recommendation:
    • NGCF (Wang et al., 2019) represents an early approach to using graph models to aggregate user and item node representations for collaborative filtering.
    • LightGCN (He et al., 2020) is a simplified and effective graph convolutional network model that achieves high performance.
    • XSimGCL (Yu et al., 2024) adopts contrastive learning to enhance representation learning on LightGCN.
  • Cross-domain Recommendation:
    • CMF (Singh and Gordon, 2008) is a classic cross-domain recommendation algorithm that employs a weighted loss to enable MF to jointly learn from both the source and target domains.
    • EMCDR (Man et al., 2017) introduces an embedding and mapping framework that involves training a mapping function to transfer source user embeddings into the target domain.
    • DeepAPF (Yan et al., 2019) leverages the attention network to capture both shared interests across sites and distinct interests within each site, capturing a diverse range of user preferences.
    • SSCDR (Kang et al., 2019) introduces a cross-domain mapping to transfer user-item distance information based on overlapping users across domains to address the cold-start issue for new users.
    • BiTGCF (Liu et al., 2020) improves the feature propagation to aggregate high-order connectivity in user-item graphs and transfer the knowledge by using the overlapping user as the bridge.
    • CUT (Li et al., 2024a) utilizes the target domain user similarity to filter collaborative information from the source domain during knowledge transfer across domains.
A.3. Additional Case Study
According to the visualization analysis in Section 4.6, we observe that the item embeddings learned by HAGO interweave across different domains, as shown in Figure 4(b). We aim to delve into the specific cross-domain correlations learned by HAGO through case study analysis. Our discussion primarily revolves around the Douban-Movie and Douban-Book domains of the Douban dataset, as the semantic relevance between items in these two domains is more intuitive and easily understood by humans.
Specifically, we select the top-5 popular books that appear in the Douban-Book domain. For each selected book, we calculate its cosine similarity with all movies in the Douban-Movie domain based on the embeddings learned by HAGO, and select the corresponding top-3 relevant movies. The resulting top-5 popular books and their corresponding top-3 relevant movies for each book are presented in Table A.3.
Table 6. Case Study: Top-5 popular books in the Douban dataset, along with top-3 relevant movies for each book thought by HAGO.
Douban-Book Douban-Movie
Item ID Item Name Item Tag Item ID Item Name Item Tag
Japanese Literature, Romance, Novel, Japan, Youth, Foreign Literature, Literature, Coming of Age. 1292220 Love Letter Unrequited Love, Aesthetic, Drama, First Love, Classic, Touching, Youth, Warmth, Romance, Literary, Japan.
1046265 Norwegian Wood 1293318 Grave of the Fireflies Family, Japanese Animation, Anime, Drama, Heartwarming, Japan, War, Animation.
1291560 My Neighbor Totoro Family, Childhood, Classic, Heartwarming, Growth, Healing, Warmth, Adventure, Fantasy, Fairy Tale, Japan, Animation.
Fairy Tales, Classic, Coming of Age, Philosophy, Foreign Literature, France, Literature, Novel. 1293839 Roman Holiday Drama, American Film, Romance, Love, America, Artistic, Film, Comedy.
1084336 The Little Prince /Le Petit Prince 1291546 Farewell My Concubine Cultural Revolution, Drama, Classic, Life, History, Romance, Humanity, Homosexuality, Literary, Artistic.
1292722 Titanic Touching, Drama, Romantic, Classic, Heartwarming, Powerful, Blockbuster, American, Disaster, Love.
Chinese Literature, Classic, Novel, Marriage, Modern Literature, Literature, Life, Romance. 1293964 Waterloo Bridge Drama, Classic, American Movies, War, American, Black and White, Romance, Movies.
1008145 Fortress Besieged /Wéichéng 1301753 The Lion King Inspirational, Family, Childhood, Classic, Coming of Age, American, Warmth, Adventure, Song and Dance, Animation.
1292271 The Graduate Drama, Classic, Life, Coming of Age, American, Love, Humanity, Youth, Comedy, Ethics.
The Unbearable Lightness of Being /Nesnesitelná lehkost bytí Foreign Literature, Novel, Philosophy, Czech, Literature, Life. 1293071 Wild Strawberries /Smultronstället European Movie, Drama, Family, Classic, Black and White, Romance, Europe, Sweden.
1017143 1439205 Panda Kopanda Japanese Animation, Ghibli, Japan, Animation.
1299080 Rear Window Crime, Thriller, Classic, American, Suspense, American Movies.
Mystery, Detective, Religious, Novel, Foreign Literature, American. 1916936 Death Note Japan, animation, anime, deduction, suspense, Japanese animation, Japanese anime, Cartoon.
1040771 The Da Vinci Code 1292262 Once Upon a Time in America Crime, Drama, Gangster, American Film, Classic, Life, Coming of Age, American.
1293544 The Silence of the Lambs Crime, Thriller, Drama, American Film, Classic, Human Nature, Psychology, American, Mystery, Horror.
For a fair comparison with other state-of-the-art baseline methods, we employ the most typical ID-based recommendation paradigm (i.e., only use user and item IDs), without using other features such as item names and tags beyond our research focus. However, as observed from Table A.3, HAGO captures many intriguing associations without the use of these semantic features: For example, the movies most closely associated with the novel “Norwegian Wood” are all Japanese films with emotional themes; those related to “The Little Prince” are some of the most classic Chinese and international films talking about love; and those linked to the novel “The Da Vinci Code” are all crime or mystery movies.
It demonstrates that cross-domain knowledge captured through multi-domain collaborative modeling of user-item interaction patterns in HAGO aligns with real semantic relevance.
A.4. Source Code
This paper is currently under review in a conference, we have submitted the source code to the venue. The source link contains the venue information so we will release it after acceptance.

### v2

@@
-Adaptive Coordinators and Prompts on Heterogeneous Graphs for Cross-Domain Recommendations
+Adaptive Graph Integration for Cross-Domain Recommendation via Heterogeneous Graph Coordinators
 Abstract.
-In the online digital world, users frequently engage with diverse items across multiple domains (e.g., e-commerce platforms, streaming services, and social media networks), forming complex heterogeneous interaction graphs.
-Leveraging this multi-domain information can undoubtedly enhance the performance of recommendation systems by providing more comprehensive user insights and alleviating data sparsity in individual domains.
-However, integrating multi-domain knowledge for the cross-domain recommendation is very hard due to inherent disparities in user behavior and item characteristics and the risk of negative transfer, where irrelevant or conflicting information from the source domains adversely impacts the target domain’s performance.
-To address these challenges, we offer HAGO, a novel framework with Heterogeneous Adaptive Graph coOrdinators, which dynamically integrate multi-domain graphs into a cohesive structure by adaptively adjusting the connections between coordinators and multi-domain graph nodes, thereby enhancing beneficial inter-domain interactions while mitigating negative transfer effects.
-Additionally, we develop a universal multi-domain graph pre-training strategy alongside HAGO to collaboratively learn high-quality node representations across domains.
-To effectively transfer the learned multi-domain knowledge to the target domain, we design an effective graph prompting method, which incorporates pre-trained embeddings with learnable prompts for the recommendation task.
-Our framework is compatible with various graph-based models and pre-training techniques, demonstrating broad applicability and effectiveness.
-Further experimental results show that our solutions outperform state-of-the-art methods in multi-domain recommendation scenarios and highlight their potential for real-world applications.
-We will open our source code [^1] This paper is currently under review in a conference, we have submitted the source code to the venue.
-The source link contains the venue information so we will release it after acceptance..
+In the digital era, users typically interact with diverse items across multiple domains (e.g., e-commerce, streaming platforms, and social networks), generating intricate heterogeneous interaction graphs.
+Leveraging multi-domain data can improve recommendation systems by enriching user insights and mitigating data sparsity in individual domains.
+However, integrating such multi-domain knowledge for cross-domain recommendation remains challenging due to inherent disparities in user behavior and item characteristics and the risk of negative transfer, where irrelevant or conflicting information from the source domains adversely impacts the target domain’s performance.
+To tackle these challenges, we propose HAGO, a novel framework with Heterogeneous Adaptive Graph coOrdinators, which dynamically integrates multi-domain graphs into a cohesive structure.
+HAGO adaptively adjusts the connections between coordinators and multi-domain graph nodes to enhance beneficial inter-domain interactions while alleviating negative transfer.
+Furthermore, we introduce a universal multi-domain graph pre-training strategy alongside HAGO to collaboratively learn high-quality node representations across domains.
+Being compatible with various graph-based models and pre-training techniques, HAGO demonstrates broad applicability and effectiveness.
+Extensive experiments show that our framework outperforms state-of-the-art methods in cross-domain recommendation scenarios, underscoring its potential for real-world applications.
+The source code is available at https://github.com/zhy99426/HAGO.
 Cross-Domain Recommendation, Graph Prompting, Graph Neural Networks
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06journalyear: 2025copyright: rightsretainedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italy.booktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italyisbn: 979-8-4007-1592-1/25/07doi: 10.1145/3726302.3729886ccs: Information systems Recommender systems
 1. Introduction
+In real-world applications, users often interact with various items (Li et al., 2024b; Meng et al., 2023b) across multiple domains within the platform, such as different product categories in e-commerce or distinct content types in streaming services, forming complex heterogeneous interaction graphs.
+However, the distribution of data across these domains is often uneven (Meng et al., 2023a).
+Some domains benefit from abundant user interaction data, while others suffer from data sparsity.
+How to effectively utilize the abundant interaction data from well-established source domains to enhance recommendation performance in the target domain that faces data sparsity or holds greater strategic importance to the platform (a.k.a, cross-domain recommendation) has become an important area in academia and industry.
 Figure 1. The schematic diagram of the existing cross-domain recommendation methods and our perspective.
-In real-world applications, users often interact with various items and other users (Li et al., 2024b) across multiple domains within the platform, such as different product categories in e-commerce or distinct content types in streaming services, forming complex heterogeneous interaction graphs.
-However, the distribution of data across these domains is often uneven.
-Some domains benefit from abundant user interaction data, while others suffer from data sparsity.
-How to effectively utilize the abundant interaction data from well-established source domains to enhance recommendation performance in the target domain that faces data sparsity or holds greater strategic importance to the platform (a.k.a cross-domain recommendation) has become an important area in academia and enterprise.
-In recent years, various model-centric methods have been proposed to handle this problem yet still leave a huge gap for further improvement.
+In recent years, various model-centric methods have been proposed to address this problem, yet significant room for improvement remains.
 As illustrated in Figure 1(a), one kind of solution is to build some shared-bottom architectures by utilizing shared representations (Singh and Gordon, 2008; Hu et al., 2018) or representation fusion (Zhu et al., 2019; Gao et al., 2019) to represent and transfer information across different domains.
@@
 Aligned with these two kinds of model-centric frameworks, they also introduce some advanced training skills such as contrastive learning, to promote cross-domain transfer learning further (Li et al., 2024a; Hao et al., 2024; Xie et al., 2022; Liu et al., 2024).
-Although these methods have achieved notable success, some later theoretical studies (Cao et al., 2023) have found that the performance ceiling of model-centric methods is strictly limited by the intrinsic discrepancy among various domains, indicating that we might not further improve these methods unless we can find some data-level approaches to narrow down the natural gap among various graph domains (as presented in Figure 1(c)).
+Although these methods have achieved notable success, some recent theoretical studies (Cao et al., 2023) have found that the performance ceiling of model-centric methods is strictly limited by the intrinsic discrepancy among various domains, indicating that we might not further improve these methods unless we can find some data-level approaches to narrow down the natural gap among various graph domains (as presented in Figure 1(c)).
 However, achieving this goal is never easy because we need to overcome at least three tough challenges: Challenge 1, designing a hand-crafted strategy to align different graph domains is far-fetched and tough.
@@
 This makes manually designed strategies far from effective and general.
-To this end, we need some trainable approaches to learn the potential latent data operation strategy to align these domains yet preserve their important knowledge.
+To this end, we need some trainable approaches to learn the potential latent data operation strategy to align these domains while preserving their important knowledge.
 Challenge 2, how to integrate and learn the rich knowledge from multiple source domains is tough.
@@
 Source domains encompass a vast array of information, much of which may not be pertinent or valuable to the target domain.
-Indiscriminately transferring knowledge from source domains to the target domain may adversely affect the model’s performance, potentially resulting in a decline in recommendation accuracy.
-Inspired by the recent success of graph prompting in its powerful data operation capability (Zhao et al., 2024; Sun et al., 2023a, c), we go beyond the previous model-centric paradigm and hope to bridge gaps between diverse graph datasets in a data-centric perspective.
-We offer HAGO framework, a data-level framework with Heterogeneous Adaptive Graph coOrdinators to address the aforementioned challenges.
-Specifically, to address Challenge 1, we revisit the theoretical basis of graph prompts in data manipulations, and then we propose graph coordinators to integrate multi-domain graphs in the pre-training stage, which shares a similar theoretical basis as graph prompts.
+Indiscriminately transferring knowledge from source domains to the target domain may adversely affect the model’s performance, potentially resulting in an accuracy decline.
+Inspired by the recent success of graph prompting in its powerful data operation capability (Zhao et al., 2024; Sun et al., 2023a, b), we go beyond the previous model-centric paradigm and hope to bridge gaps between diverse graph datasets in a data-centric perspective.
+We offer HAGO framework, a data-level framework with Heterogeneous Adaptive Graph coOrdinators to address the above challenges.
+Specifically, to address Challenge 1, we revisit the theoretical basis of graph prompts in data manipulations, and then we propose heterogeneous graph coordinators to integrate multi-domain graphs in the pre-training stage, which shares a similar theoretical basis as graph prompts.
 In this way, we can get rid of hand-crafted strategies and learn a potential way to unify various source domains into a unified structure, enabling collaborative multi-domain pre-training to facilitate representation alignment among them.
 Notably, our graph coordinators are designed to be heterogeneous, using user/item coordinators to connect user/item nodes in graphs, thus preventing negative transfer caused by information confusion between different types of nodes.
-To address Challenge 2, we give the details of a multi-domain pre-training framework with our graph coordinator mechanism.
+To address Challenge 2, we present the details of a multi-domain pre-training framework with our graph coordinator mechanism.
 This framework is model-agnostic and can be easily used in various pre-training strategies and backbones.
@@
 2. Background
-In this section, we introduce our motivation by comparing two perspectives of cross-domain recommendation.
-Then, the formal objective of this paper will be given.
-We give a more comprehensive review of existing related work in Appendix A.1.
+In this section, we introduce our motivation by comparing two perspectives of cross-domain recommendation. Then, the formal objective of this paper will be given.
 Cross-domain Recommendation: A Model-centric Perspective.
-A mainstream approach for cross-domain recommendation is to employ a shared-bottom architectures (Singh and Gordon, 2008; Hu et al., 2018; Zhu et al., 2019; Gao et al., 2019; Zhu et al., 2024) and mapping functions to project knowledge from source domains to the target (Man et al., 2017; Zhao et al., 2020; Kang et al., 2019; Guo et al., 2023).
+A mainstream approach for cross-domain recommendation is to employ a shared-bottom architecture (Singh and Gordon, 2008; Hu et al., 2018; Zhu et al., 2019; Gao et al., 2019; Zhu et al., 2024) and mapping functions to project knowledge from source domains to the target (Man et al., 2017; Zhao et al., 2020; Kang et al., 2019; Guo et al., 2023).
 These approaches are inherently model-centric.
-Although much progress achieved, some researchers recently have found that model-centric approaches have an unbreakable ceiling caused by the intrinsic gap of the cross-domain data (Cao et al., 2023).
-That means no matter how better the models are, their upper bound performance is always limited by the data natural gap.
-Thus, it is imperative to return to the intrinsic nature of the problem and consider how to further break up the ceiling.
-Intuitively, if we can find some way to narrow the data gap among various domains, we will have a higher performance improvement margin for existing models.
-However, traditional data dealing approaches heavily rely on manpower and it is far from effective and efficient to manually design data operation strategies for this vision.
-Figure 2.
-Illustration of the proposed HAGO, using two source domains and a target domain as examples.
-There is partial overlap among user nodes across different domains.
-For the sake of clarity in the diagram, we omit the overlapping user nodes that are present in all domains.
-HAGO employs heterogeneous adaptive graph coordinators to organize the multi-domain pre-training collaboratively, and then leverages graph prompts to transfer the knowledge gained from pre-training to the target domain, thereby enhancing downstream recommendation tasks.
+Despite significant progress, recent studies have revealed that model-centric approaches have an unbreakable ceiling caused by the intrinsic gap of the cross-domain data (Cao et al., 2023).
+That means no matter how good the models are, their upper bound performance is always limited by the natural data gap.
+Thus, it is imperative to revisit the intrinsic nature of the problem and explore ways to further break up the ceiling.
+Intuitively, narrowing the data gap among various domains could provide a higher performance improvement margin for existing models.
+However, traditional data dealing approaches heavily rely on manpower, making it inefficient and ineffective to manually design data operation strategies for this vision.
 Our Motivation: An Attempt of Data-centric Perspective.
 With the above discussion, this paper tries to explore an enhanced direction, which aims to achieve representation alignment and effective knowledge transfer from a data-centric perspective.
-Unlike other linear-structured data such as images and text, graph data is highly complex and possesses diverse non-linear structural properties.
-The nodes in a graph are not isolated; they are deeply influenced by their local neighbors.
+Unlike other linear-structured data such as images and text, graph data is highly complex and exhibits diverse non-linear structural properties.
+Nodes in a graph are not isolated; they are deeply influenced by their local neighbors.
 The connectivity among nodes serves as the carrier of information flow within the graph, reflecting the core characteristics of graph data.
-Consequently, capturing the common patterns of complex and diverse local structures across multiple graphs to align graph data at the data level particularly presents a more challenging problem.
+Consequently, capturing the common patterns (Zhang et al., 2024) of complex and diverse local structures across multiple graphs to align graph data at the data level particularly presents a more challenging problem.
 It is infeasible to leverage manually designed strategies to heal the dataset gap effectively.
-Luckily, recent progress on graph prompting has demonstrated the potential effectiveness in manipulating diversified datasets to be unified (Wang et al., 2024).
+Fortunately, recent progress on graph prompting has demonstrated the potential effectiveness in manipulating diversified datasets to be unified (Sun et al., 2023a; Wang et al., 2024).
 Graph prompting has the potential to bridge the gaps between multi-domain pre-training and knowledge transfer.
-With this in mind, we aspire to harness the power of graph prompting to amalgamate graphs from various domains into a cohesive structure, fostering alignment and mitigating negative transfer at the data level.
+Inspired by this, we aspire to harness the power of graph prompting to amalgamate graphs from various domains into a cohesive structure, fostering alignment and mitigating negative transfer at the data level.
 Formalized Objective.
@@
 The objective is to sufficiently utilize the rich information from source domains $\{\mathcal{G}^{(1)},\cdots,\mathcal{G}^{(k)},\cdots,\mathcal{G}^{(M)}\}$ to improve the Top-N recommendation performance in the target domain $\mathcal{G}^{(t)}=(\mathcal{V}^{(t)},\mathcal{E}^{(t)})$ with the adjacency matrix $\boldsymbol{A}^{(t)}$.
-3. Reconcile Multi-Domain Incompatibility via Graph Coordinators and Prompts
+3. Method
+Figure 2.
+Illustration of the proposed HAGO, using two source domains and a target domain as examples.
+There is partial overlap among user nodes across different domains.
+For clarity, we omit the overlapping user nodes that are present in all domains.
+HAGO employs heterogeneous adaptive graph coordinators to organize the multi-domain pre-training collaboratively, and then leverages graph prompts to transfer the knowledge gained from pre-training to the target domain, thereby enhancing downstream recommendation tasks.
 Overview: With the above objective, we can see that we should not only try to narrow the gap between the source domains and the target domain, but also need to heal the incompatibility among multiple source domains.
-To this end, we proposed HAGO as depicted in Figure 2.
+To this end, the proposed HAGO is depicted in Figure 2.
 Specifically, our framework initially employs heterogeneous adaptive graph coordinators to integrate multi-domain graphs into a unified structure and leverages self-supervised learning algorithms for collaborative multi-domain pre-training.
@@
 (1) $\boldsymbol{e}_{u}=\boldsymbol{E}^{T}_{u}\boldsymbol{o}_{u},\boldsymbol{e}_{i}=\boldsymbol{E}^{T}_{i}\boldsymbol{o}_{i},$
-where $\boldsymbol{E}_{u}\in\mathbb{R}^{|\mathcal{U}|\times d}$ and $\boldsymbol{E}_{i}\in\mathbb{R}^{|\mathcal{I}|\times d}$ denotes the embedding lookup table of users and items, and $\boldsymbol{o}_{u}\in\mathbb{R}^{|\mathcal{U}|},\boldsymbol{o}_{i}\in\mathbb{R}^{|\mathcal{I}|}$ are the corresponding one-hot vector of user/item IDs.
+where $\boldsymbol{E}_{u}\in\mathbb{R}^{|\mathcal{U}|\times d}$ and $\boldsymbol{E}_{i}\in\mathbb{R}^{|\mathcal{I}|\times d}$ denote the user/item embedding lookup tables, and $\boldsymbol{o}_{u}\in\mathbb{R}^{|\mathcal{U}|},\boldsymbol{o}_{i}\in\mathbb{R}^{|\mathcal{I}|}$ are the corresponding one-hot vector of user/item IDs.
 $|\mathcal{U}|$ and $|\mathcal{I}|$ indicate the numbers of users and items, and $d$ means the embedding size.
-Propagation Layer: The graph aggregation process of LightGCN can be formalized as:
-(2) $\boldsymbol{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}\boldsymbol{e}_{i}^{(l)},$
-$\boldsymbol{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{u}\right|}}\boldsymbol{e}_{u}^{(l)},$
+Propagation Layer: The graph aggregation process in LightGCN can be formalized as:
+(2) $\boldsymbol{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}\boldsymbol{e}_{i}^{(l)},\boldsymbol{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{u}\right|}}\boldsymbol{e}_{u}^{(l)},$
 where $\boldsymbol{e}_{u}^{(l)}$ and $\boldsymbol{e}_{i}^{(l)}$ mean the latent representations of user $u$ and item $i$ after $l$-layer graph convolutions, and $\boldsymbol{e}^{(0)}_{u}=\boldsymbol{e}_{u}$, $\boldsymbol{e}^{(0)}_{i}=\boldsymbol{e}_{i}$.
@@
 3.2. Heterogeneous Adaptive Graph Coordinator
+To enhance the interconnections between multi-domains while alleviating the negative transfer phenomenon, we design effective heterogeneous adaptive graph coordinators to organize cross-domain knowledge transfer dynamically.
 3.2.1. Heterogeneous Graph Coordinators
@@
 Naturally, we introduce a set of user coordinators $\mathcal{C}_{u}^{(k)}=\{c_{u,1}^{(k)},c_{u,2}^{(k)},\cdots,c_{u,n}^{(k)}\}$ and item coordinators $\mathcal{C}_{i}^{(k)}=\{c_{i,1}^{(k)},c_{i,2}^{(k)},\cdots,$ $c_{i,n}^{(k)}\}$ for graph $\mathcal{G}^{(k)}$.
-The whole set of graph coordinators is $\mathcal{C}=\bigcup_{k}\mathcal{C}_{u}^{(k)}\cup\bigcup_{k}\mathcal{C}_{i}^{(k)}$, and the number of coordinators is $|\mathcal{C}|=2n$.
+Here, user coordinators are restricted to connecting only with user nodes, while item coordinators are restricted to connecting only with item nodes.
+The hyperparameter $n$ determines the number of coordinators for each type, meaning we introduce $2n$ coordinators including $n$ user coordinators and $n$ item coordinators.
 Each graph coordinator connects to the nodes in the corresponding graph through dynamically weighted direct edges, and corresponds to a learnable embedding which can be obtained by retrieving the coordinator embedding table $\boldsymbol{E}_{c}\in\mathbb{R}^{|\mathcal{C}|\times d}$ as the same way in Equation 1.
@@
 To enhance the effectiveness of cross-domain knowledge transfer while alleviating the negative transfer phenomenon, we propose an adaptive edge weight learning strategy to dynamically regulate the strength of information transfer between the coordinators and the nodes.
-Our proposed strategy encompasses both soft/hard cases and can be formalized as follows:
+Our proposed strategy includes both soft/hard cases and can be formalized as follows:
 (3) $w=\left\{\begin{aligned} \frac{\boldsymbol{e}\cdot\boldsymbol{e}_{c}}{||\boldsymbol{e}||||\boldsymbol{e}_{c}||},&\ \ \textit{if}\ \ \boldsymbol{e}\cdot\boldsymbol{e}_{c}>0,\qquad\textit{(Soft Case)}\\ 0\qquad,&\ \ \textit{if}\ \ \boldsymbol{e}\cdot\boldsymbol{e}_{c}\leq 0,\qquad\textit{(Hard Case)}\end{aligned}\right.$
@@
 3.2.3. Interconnections between Coordinators
-Through the aforementioned strategy, we establish adaptive linkages between graph coordinators and their respective graphs, empowering these coordinators to bolster individual graph models by regulating the flow of information among the nodes.
-Our vision, however, transcends this capability, as we aspire to model collaborative connections across multiple domains within our dataset to facilitate multi-domain representation alignment.
+Through the above strategy, we establish adaptive linkages between graph coordinators and their respective graphs, empowering these coordinators to bolster individual graph models by regulating the flow of information among the nodes.
+Our vision extends further, aiming to model collaborative connections across multiple domains to facilitate multi-domain representation alignment.
 Consequently, we establish direct cross-domain interconnections between the graph coordinators associated with each domain graph, effectively amalgamating the entire collection of graphs into a cohesive structure.
 These graph coordinators assume the pivotal role of information exchange hub between graphs, with the connecting edges between coordinators acting as bridges for this communication.
-Notably, our proposed HAGO framework includes distinct coordinators for users and items, facilitating connections solely between coordinators of different types.
-This structured relationship can be effectively depicted using an adjacency matrix $\boldsymbol{R}_{C}$:
+Notably, our proposed HAGO framework includes distinct coordinators for users and items, facilitating connections only between coordinators of different types.
+In other words, user-user and item-item coordinator connections are not included.
+This design is based on the intuition that, even with the introduction of graph coordinators, the cohesive multi-domain graph should retain the structure of a heterogeneous user-item bipartite graph.
+This structured relationship can be depicted using an adjacency matrix $\boldsymbol{R}_{C}$:
 (4) $\begin{array}[]{cc}\boldsymbol{R}_{C}=\left[\begin{array}[]{cc}\boldsymbol{0}_{|\mathcal{C}_{u}|\times|\mathcal{C}_{u}|}&\boldsymbol{J}_{|\mathcal{C}_{u}|\times|\mathcal{C}_{i}|}\\ \boldsymbol{J}_{|\mathcal{C}_{i}|\times|\mathcal{C}_{u}|}&\boldsymbol{0}_{|\mathcal{C}_{i}|\times|\mathcal{C}_{i}|}\end{array}\right],&\begin{array}[]{cc}\leftarrow user\\ \leftarrow item\end{array}\\ \begin{array}[]{cc}\uparrow&\qquad\uparrow\\ user&\qquad item\end{array}&\end{array}$
@@
 GRACE (Zhu et al., 2020) uses graph augmentation techniques such as edge removing and feature masking to generate pairs of positive samples, and then it employs the InfoNCE (van den Oord et al., 2018) loss function to ensure that these positive samples are brought close together while negative samples are pushed apart.
-In order to demonstrate the wide versatility of our framework, we subsequently adopt other state-of-the-art graph pre-training methods like BGRL and COSTA, integrating them into our multi-domain pre-training framework, in Section 4.4.
+In order to demonstrate the wide versatility of our framework, we subsequently adopt other state-of-the-art graph pre-training methods like BGRL (Thakoor et al., 2021) and COSTA (Zhang et al., 2022), integrating them into our multi-domain pre-training framework, in Section 4.4.
 COSTA builds upon GRACE by incorporating the novel Covariance-Preserving Feature Augmentation technique, which effectively tackles the bias challenges present in the original augmentation strategies.
@@
 $=-\frac{1}{|\mathcal{B}|}\sum\limits_{(u,i,j)\in\mathcal{B}}\mathrm{ln}\sigma(\hat{y}_{ui}-\hat{y}_{uj})+\lambda||\Theta||^{2},$
-where the second term is the $2$-regularization for potential overfitting issues.
+where the second term is the $L2$-regularization to prevent overfitting.
 $(u,i,j)$ from batch $\mathcal{B}$ is a triple including user $u$ and item $i,j$.
@@
 Theoretical Support: Our proposed framework shares a common theoretical bedrock with graph prompts designed for knowledge transfer.
-The primary difference is that while graph prompts are generally targeted at downstream tasks, we utilize the term ”graph coordinators” for the pre-training phase.
-Studies (Wang et al., 2024) have shown that for a given graph $\mathcal{G}$, we can find some potential graph transformation $T(\cdot)$ to translate the original graph as a new one like $T(\mathcal{G})$ such that the pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ can approximate to an optimal function $C$ for the target domain task, which can be described as:
+We draw inspiration from the “Prompt as Graph” paradigm (Zi et al., 2024; Sun et al., 2023b), which introduces multiple additional prompt nodes with the inner structure, which correspond to learnable embeddings, into the original graph and connects them through specific structures (Sun et al., 2023a; Huang et al., 2023).
+The primary difference is that while graph prompts are generally targeted at downstream tasks, we utilize the term “graph coordinators” for the pre-training phase.
+Previous studies (Fang et al., 2023; Sun et al., 2023a) have demonstrated that graph prompts can emulate any graph operations, such as node/edge addition and deletion.
+Recent studies (Wang et al., 2024) have shown that for a given graph $\mathcal{G}$, we can find some potential graph transformation $T(\cdot)$ to translate the original graph as a new one like $T(\mathcal{G})$ such that the pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ can approximate to an optimal function $C$ for the target domain task, which can be described as:
 (8) $f_{\theta^{*}}(T(\mathcal{G}))=C(\mathcal{G}).$
-It can be proven that graph prompts can be treated as a learnable method to find such data transformation (Fang et al., 2023; Sun et al., 2023a).
+In other words, graph prompts can be treated as a learnable method to find such data transformation (Fang et al., 2023; Sun et al., 2023a) to approximate optimal in the target task (Wang et al., 2024).
 Specifically, given a pre-trained graph encoder $f_{\theta^{*}}(\cdot)$ with fixed parameters $\theta^{*}$ and learnable graph prompt $P(\cdot)$, it can be shown that:
@@
 where $O_{fp}$ is the bounded error associated with $f_{\theta^{*}}(\cdot)$ and $P(\cdot)$.
-In other words, graph prompts have the capability to be equivalent to certain graph transformation operations that enable the transformed graph, once processed by the pre-trained graph encoder, to produce optimal embeddings.
-In this paper, our heterogeneous adaptive graph coordinators can be treated as a variant of graph prompts, which is used in the pre-training stage, and thus share the same foundation for data operations.
+This implies that graph prompts have the capability to be equivalent to certain graph transformation operations that enable the transformed graph, once processed by the pre-trained graph encoder, to produce optimal embeddings.
+In this paper, our heterogeneous adaptive graph coordinators can be treated as a variant of graph prompts, which are used in the pre-training stage, and thus share the same foundation for data operations.
 Based on this idea, we give a data-centric solution to heal the gap among multiple source domains and the target domain such that the performance ceiling of knowledge transfer can be further improved as mentioned in our motivation (Section 2).
 Compatibility Analysis: Our proposed HAGO is model-agnostic because it extends the graph structure without regarding to message passing and aggregation within the graph, making it compatible with various graph-based backbone networks.
-Graph Prompting can also be seamlessly integrated with various graph pre-training algorithms.
-In our subsequent experiments, comprehensive compatibility analysis demonstrates the broad compatibility and wide effectiveness of HAGO with different backbone networks and pre-training methods.
+It can also be seamlessly integrated with various graph pre-training algorithms.
+In our subsequent experiments, comprehensive compatibility analysis demonstrates the broad compatibility and wide effectiveness of HAGO with different backbones and pre-training methods.
 Complexity Analysis: The time complexity of our proposed HAGO framework is intrinsically linked to that of the backbone model.
@@
 Consequently, for HAGO, the time complexity during the multi-domain pre-training phase is $\mathcal{O}((\sum\limits_{k=1}^{M}|\mathcal{E}^{(k)}|+|\mathcal{E}^{(t)}|+|\mathcal{E}_{C}|)Ld)$ where $|\mathcal{E}_{C}|$ is the number of edges introduced by HAGO; and when transferring to downstream recommendation tasks via graph prompts, the time complexity remains $\mathcal{O}(|\mathcal{E}^{(t)}|Ld)$.
-4. Evaluation Analysis
-4.1. Experiment Settings
-Table 1. Dataset Statistics
-dataset domain #users #items #interactions sparsity
-Douban Music 15,996 39,749 1,116,984 99.82%
-Movie 22,041 25,802 2,552,305 99.55%
-Book 18,086 33,068 809,248 99.86%
-Amazon Book 135,110 115,173 4,042,382 99.97%
-Electronic 112,988 41,905 1,880,865 99.96%
-Movie 26,969 18,564 762,957 99.85%
-Toy 11,687 8,289 202,230 99.79%
 Table 1.
@@
 Amazon MRR 0.0601 0.1074 0.1088 0.0542 0.0752 0.0951 0.0963 0.1056 0.1055 0.1116
-Data Preparation and Setup: We conduct comprehensive experiments on two popular public recommendation datasets: Amazon dataset (Ni et al., 2019) [^2] http://jmcauley.ucsd.edu/data/amazon is gathered from relevant product evaluations on Amazon.com, which are extensively utilized to assess the performance of cross-domain recommender systems.
+4. Evaluation Analysis
+4.1. Experiment Settings
+Table 2. Dataset Statistics
+dataset domain #users #items #interactions sparsity
+Douban Music 15,996 39,749 1,116,984 99.82%
+Movie 22,041 25,802 2,552,305 99.55%
+Book 18,086 33,068 809,248 99.86%
+Amazon Book 135,110 115,173 4,042,382 99.97%
+Electronic 112,988 41,905 1,880,865 99.96%
+Movie 26,969 18,564 762,957 99.85%
+Toy 11,687 8,289 202,230 99.79%
+Data Preparation and Setup: We conduct comprehensive experiments on two popular public datasets: Amazon dataset (Ni et al., 2019) [^1] http://jmcauley.ucsd.edu/data/amazon is gathered from relevant product evaluations on Amazon.com.
 For our study, we have chosen four categories: “Books”, “Electronics”, “Movies and TV”, and “Toys and Games”; Douban dataset (Zhao et al., 2021) [^2] https://recbole.s3-accelerate.amazonaws.com/CrossDomain/Douban.zip is collected from Douban website, which is a large-scale platform for rating and reviewing movies, TV shows, books, and music.
@@
 Evaluation Metrics: We assess the performance of our method by four widely used metrics in recommendations: Recall@10, HR@10 (Hit Ratio), NDCG@10 (Normalized Discounted Cumulative Gain), and MRR (Mean Reciprocal Rank) for top-10 recommendation.
-Kindly note that we fully rank the entire items within the datasets for calculating the above top-10 recommendation performance rather than evaluate these metrics by pre-sampling a small group of the total items.
+Kindly note that we fully rank the entire items within the datasets for calculating the above top-10 recommendation performance rather than evaluating these metrics by pre-sampling a small group of the total items.
 This is closer to the real industrial applications and raises more challenging tasks for the compared models.
-Baselines: We compare our method with NINE state-of-the-art methods, which can be grouped into two categories: (1) Single-domain Recommendation Methods withNGCF (Wang et al., 2019), LightGCN (He et al., 2020), and XSimGCL (Yu et al., 2024).
+Baselines: We compare our method with NINE state-of-the-art (SOTA) methods, which can be grouped into two categories: (1) Single-domain Recommendation Methods with NGCF (Wang et al., 2019), LightGCN (He et al., 2020), and XSimGCL (Yu et al., 2024).
 (2) Cross-domain Recommendation Methods with CMF (Singh and Gordon, 2008), EMCDR (Man et al., 2017), DeepAPF (Yan et al., 2019), SSCDR (Kang et al., 2019), BiTGCF (Liu et al., 2020), and CUT (Li et al., 2024a).
-The details of these methods are present in Appendix A.2.
 Implementation Details: All the baselines and our proposed HAGO are implemented based on the PyTorch (Paszke et al., 2019).
@@
   • Our proposed HAGO consistently achieves the best performance across all four metrics on two multi-domain datasets, demonstrating its superior effectiveness.
-It gains 1.20% and 1.28% improvements on Recall@10, 1.44% and 1.67% on HR@10, 1.32% and 1.64% on NDCG@10, and 0.59% and 2.57% on MRR against the runner-up on Douban and Amazon datasets.
 HAGO owes its success to employing graph coordinators that seamlessly integrate multi-domain graphs into a cohesive framework for multi-domain collaborative pre-training.
@@
 4.4. Compatibility Analysis with Different Backbones and Pre-training Strategies
+Our proposed HAGO is not limited to a specific model but is designed as a flexible architecture that can be effortlessly combined with a variety of backbone networks serving as graph encoders, as well as with diverse graph pre-training algorithms.
+To demonstrate the adaptability of HAGO, we conduct comparative experiments by integrating it with other notable graph neural networks, including GCN (Kipf and Welling, 2017) and GAT (Velickovic et al., 2018).
+Furthermore, we investigate the incorporation of other self-supervised learning strategies for pre-training the graph, such as BGRL (Thakoor et al., 2021) and COSTA (Zhang et al., 2022).
 Table 4.
@@
 HAGO(GAT) 0.1248 0.2416 0.0850 0.1043
-LightGCN 0.1665 0.3116 0.1210 0.1516
-HAGO(LightGCN) 0.1689 0.3168 0.1227 0.1535
 Amazon
@@
 HAGO(GAT) 0.1350 0.1571 0.0939 0.0874
-LightGCN 0.1627 0.1902 0.1141 0.1074
-HAGO(LightGCN) 0.1662 0.1950 0.1177 0.1116
-Our proposed HAGO is not confined to a specific model but is designed as a flexible architecture that can be effortlessly combined with a variety of backbone networks serving as graph encoders, as well as with diverse graph pre-training algorithms.
-To demonstrate the adaptability of HAGO, we conduct comparative experiments by integrating it with other notable graph neural networks, including GCN and GAT.
-Furthermore, we investigate the incorporation of other self-supervised learning strategies for pre-training the graph, such as BGRL and COSTA.
 Table 5.
@@
 Figure 3. Performances with different numbers of graph coordinators $2n$.
-Observing the results, it is evident that the optimal performance of HAGO is reached when the number of graph coordinators is moderate.
+Observing the results, it is evident that the optimal performance of HAGO is reached when the number of coordinators is moderate.
 When there are too few coordinators, they are inadequate to handle the demanding task of transferring knowledge both within and across domains efficiently.
@@
 4.6. Visualization Analysis
-To investigate whether the HAGO framework we propose can effectively align the representation spaces and facilitate cross-domain knowledge transfer among multiple domains, we conduct a visualization analysis of the pre-trained embeddings for item nodes in each domain, following the analytical methods in (Yu et al., 2022).
+To investigate whether the proposed HAGO framework can effectively align the representation spaces and facilitate cross-domain knowledge transfer among multiple domains, we conduct a visualization analysis of the pre-trained embeddings for item nodes in each domain, following the analytical methods in (Yu et al., 2022).
 We randomly sample 2000 items from each of the three domains in the Douban dataset (Douban-Music, Douban-Movie, and Douban-Book).
@@
 This observation underscores the effectiveness of HAGO in enhancing the expressiveness of embeddings by promoting a more cohesive structure within multi-domain graphs, thereby facilitating knowledge transfer across domains.
-We give more additional case studies in Appendix A.3.
+4.7. Case Study
+According to the visualization analysis in Section 4.6, we observe that the item embeddings learned by HAGO interweave across different domains, as shown in Figure 4(b).
+We aim to delve into the specific cross-domain correlations learned by HAGO through case study analysis.
+Our discussion primarily revolves around the Douban-Movie and Douban-Book domains of the Douban dataset, as the semantic relevance between items in these two domains is more intuitive and easily understood by humans.
+Specifically, we select the most popular books under four specific themes (including romance, thriller, historical, and fantasy) that appear in the Douban-Book domain.
+For each selected book, we calculate its cosine similarity with all movies in the Douban-Movie domain based on the item embeddings learned by HAGO, and select the corresponding top-3 relevant movies.
+The resulting most popular books and their corresponding top-3 relevant movies for each book are presented in Figure 5.
+Figure 5. Case Study: The most popular books under four specific themes in the Douban dataset, along with top-3 relevant movies for each book thought by HAGO.
+For a fair comparison with other SOTA baseline methods, we employ the most typical ID-based recommendation paradigm (i.e., only use user and item IDs), without using other features such as item names and tags beyond our research focus.
+However, as observed from Figure 5, HAGO captures many intriguing associations without the use of these semantic features: For instance, the movies most closely associated with the romance novel “Norwegian Wood” are all Japanese films with emotional themes (like love, family, and friendship).
+In contrast, those tied to the thriller novel “The Da Vinci Code” are primarily crime or mystery genres.
+Moreover, users who enjoy the historical novel “Stories about Ming Dynasty” may also like some movies talking about World War II or Chinese historical stories, while movies associated with “Harry Potter and the Philosopher’s Stone” are all children’s favorites such as animations, Chinese Xianxia, and Disney adventures.
+It demonstrates that cross-domain knowledge captured through multi-domain collaborative modeling of user-item interaction patterns in HAGO aligns with real semantic relevance.
 5. Conclusions
@@
 To tackle these intricate challenges, we introduce the HAGO framework, designed to integrate multi-domain graphs into a cohesive structure.
-Within HAGO, graph coordinators dynamically connect with corresponding graphs based on adaptive weights, thereby regulating intra-domain and inter-domain information transfer to mitigate the negative transfer problem.
 We conduct comprehensive experiments to demonstrate the effectiveness of HAGO and vividly discuss its role in representation alignment and the facilitation of cross-domain transfer through visual analysis and case studies.
-Appendix A Appendix
-The Appendix is organized as follows: We supplement a more comprehensive review of related work in Appendix A.1, then give a more detailed introduction to our compared baselines in Appendix A.2.
-In Appendix A.3, we extend our experimental discussion with additional case studies.
-We open our source code in Appendix A.4.
-A.1. Related Work
-A.1.1. Self-Supervised Graph Learning
-Self-supervised graph learning is a machine learning algorithm that trains graph models using supervision signals generated from the graph data itself, without relying on additional labels (Sun et al., 2023b).
-Specifically, self-supervised graph learning algorithms can be categorized into generative, predictive, and contrastive approaches  (Liu et al., 2023; Wu et al., 2023).
-Generative methods typically rely on an encoder-decoder structure to capture generalized latent representations through tasks like feature or structure reconstruction, such as MGAE (Wang et al., 2017) and GALA (Park et al., 2019) for feature reconstruction, and GAE (Kipf and Welling, 2016) and GraphMAE (Hou et al., 2022, 2023) for structure reconstruction.
-Predictive methods use pseudo-labels generated from the data itself as supervision signals.
-For instance, S2GRL (Peng et al., 2020) and PairwiseDistance (Jin et al., 2020) use the prediction of the shortest path between two nodes, while M3S (Sun et al., 2020) and SHGP (Yang et al., 2022) generate pseudo-labels through clustering.
-Contrastive methods, which are the most widely studied, learn representations by constraining positive samples to be closer and pushing negative samples apart.
-These methods typically augment graph data by adding or removing edges, dropping nodes, feature transformations, or dropping features to generate a series of positive samples, then utilize loss functions such as InfoNCE  (Zhu et al., 2020; You et al., 2020; Xia et al., 2022; Zhu et al., 2022; Zhang et al., 2022), Jensen-Shannon Divergence  (Hassani and Ahmadi, 2020; Sun et al., 2021), or Bootstrapping Latent Loss  (Thakoor et al., 2021) to constrain the distance between positive and negative samples.
-A.1.2. Cross-domain Recommendation
-Cross-domain recommendation refers to the recommendation algorithms that enhance recommendations in one domain by transferring knowledge from other domains.
-These algorithms can be categorized into two types: those without overlapping users/items and those with overlapping users/ items.
-In non-overlapping cases, research mainly focuses on transferring or aligning semantic information across domains to facilitate knowledge transfer between the source and target domains  (Li et al., 2009; Liu et al., 2022).
-For overlapping cases, studies leverage shared users and/or items to transfer knowledge from the source domains to the target domain.
-Early attempts utilize shared representations (such as CMF (Singh and Gordon, 2008)) or concatenated representations from different domains (such as DTCDR (Zhu et al., 2019)) to make predictions, learning from both the source and target domains.
-Later researchers explore building cross-domain connections between neural network models to enhance knowledge transfer, such as CoNet (Hu et al., 2018) and BiTGCF (Liu et al., 2020).
-EMCDR (Man et al., 2017) is the first to propose a mapping function to transfer knowledge at the latent representation level, which is further expanded by models like CATN (Zhao et al., 2020) and SSCDR (Kang et al., 2019).
-More recent studies, such as CUT (Li et al., 2024a), have begun exploring the utilization of contrastive learning and other self-supervised techniques to transfer useful knowledge from the source to the target domain, improving recommendation accuracy.
-A.2. Details of Baselines
-In order to thoroughly assess the effectiveness of our proposed HAGO, we conduct a comprehensive comparison with nine state-of-the-art cross-domain baseline models:
-  • Single-domain Recommendation:
-    • NGCF (Wang et al., 2019) represents an early approach to using graph models to aggregate user and item node representations for collaborative filtering.
-    • LightGCN (He et al., 2020) is a simplified and effective graph convolutional network model that achieves high performance.
-    • XSimGCL (Yu et al., 2024) adopts contrastive learning to enhance representation learning on LightGCN.
-  • Cross-domain Recommendation:
-    • CMF (Singh and Gordon, 2008) is a classic cross-domain recommendation algorithm that employs a weighted loss to enable MF to jointly learn from both the source and target domains.
-    • EMCDR (Man et al., 2017) introduces an embedding and mapping framework that involves training a mapping function to transfer source user embeddings into the target domain.
-    • DeepAPF (Yan et al., 2019) leverages the attention network to capture both shared interests across sites and distinct interests within each site, capturing a diverse range of user preferences.
-    • SSCDR (Kang et al., 2019) introduces a cross-domain mapping to transfer user-item distance information based on overlapping users across domains to address the cold-start issue for new users.
-    • BiTGCF (Liu et al., 2020) improves the feature propagation to aggregate high-order connectivity in user-item graphs and transfer the knowledge by using the overlapping user as the bridge.
-    • CUT (Li et al., 2024a) utilizes the target domain user similarity to filter collaborative information from the source domain during knowledge transfer across domains.
-A.3. Additional Case Study
-According to the visualization analysis in Section 4.6, we observe that the item embeddings learned by HAGO interweave across different domains, as shown in Figure 4(b).
-We aim to delve into the specific cross-domain correlations learned by HAGO through case study analysis.
-Our discussion primarily revolves around the Douban-Movie and Douban-Book domains of the Douban dataset, as the semantic relevance between items in these two domains is more intuitive and easily understood by humans.
-Specifically, we select the top-5 popular books that appear in the Douban-Book domain.
-For each selected book, we calculate its cosine similarity with all movies in the Douban-Movie domain based on the embeddings learned by HAGO, and select the corresponding top-3 relevant movies.
-The resulting top-5 popular books and their corresponding top-3 relevant movies for each book are presented in Table A.3.
-Table 6. Case Study: Top-5 popular books in the Douban dataset, along with top-3 relevant movies for each book thought by HAGO.
-Douban-Book Douban-Movie
-Item ID Item Name Item Tag Item ID Item Name Item Tag
-Japanese Literature, Romance, Novel, Japan, Youth, Foreign Literature, Literature, Coming of Age.
-1292220 Love Letter Unrequited Love, Aesthetic, Drama, First Love, Classic, Touching, Youth, Warmth, Romance, Literary, Japan.
-1046265 Norwegian Wood 1293318 Grave of the Fireflies Family, Japanese Animation, Anime, Drama, Heartwarming, Japan, War, Animation.
-1291560 My Neighbor Totoro Family, Childhood, Classic, Heartwarming, Growth, Healing, Warmth, Adventure, Fantasy, Fairy Tale, Japan, Animation.
-Fairy Tales, Classic, Coming of Age, Philosophy, Foreign Literature, France, Literature, Novel. 1293839 Roman Holiday Drama, American Film, Romance, Love, America, Artistic, Film, Comedy.
-1084336 The Little Prince /Le Petit Prince 1291546 Farewell My Concubine Cultural Revolution, Drama, Classic, Life, History, Romance, Humanity, Homosexuality, Literary, Artistic.
-1292722 Titanic Touching, Drama, Romantic, Classic, Heartwarming, Powerful, Blockbuster, American, Disaster, Love.
-Chinese Literature, Classic, Novel, Marriage, Modern Literature, Literature, Life, Romance. 1293964 Waterloo Bridge Drama, Classic, American Movies, War, American, Black and White, Romance, Movies.
-1008145 Fortress Besieged /Wéichéng 1301753 The Lion King Inspirational, Family, Childhood, Classic, Coming of Age, American, Warmth, Adventure, Song and Dance, Animation.
-1292271 The Graduate Drama, Classic, Life, Coming of Age, American, Love, Humanity, Youth, Comedy, Ethics.
-The Unbearable Lightness of Being /Nesnesitelná lehkost bytí Foreign Literature, Novel, Philosophy, Czech, Literature, Life.
-1293071 Wild Strawberries /Smultronstället European Movie, Drama, Family, Classic, Black and White, Romance, Europe, Sweden.
-1017143 1439205 Panda Kopanda Japanese Animation, Ghibli, Japan, Animation.
-1299080 Rear Window Crime, Thriller, Classic, American, Suspense, American Movies.
-Mystery, Detective, Religious, Novel, Foreign Literature, American. 1916936 Death Note Japan, animation, anime, deduction, suspense, Japanese animation, Japanese anime, Cartoon.
-1040771 The Da Vinci Code 1292262 Once Upon a Time in America Crime, Drama, Gangster, American Film, Classic, Life, Coming of Age, American.
-1293544 The Silence of the Lambs Crime, Thriller, Drama, American Film, Classic, Human Nature, Psychology, American, Mystery, Horror.
-For a fair comparison with other state-of-the-art baseline methods, we employ the most typical ID-based recommendation paradigm (i.e., only use user and item IDs), without using other features such as item names and tags beyond our research focus.
-However, as observed from Table A.3, HAGO captures many intriguing associations without the use of these semantic features: For example, the movies most closely associated with the novel “Norwegian Wood” are all Japanese films with emotional themes; those related to “The Little Prince” are some of the most classic Chinese and international films talking about love; and those linked to the novel “The Da Vinci Code” are all crime or mystery movies.
-It demonstrates that cross-domain knowledge captured through multi-domain collaborative modeling of user-item interaction patterns in HAGO aligns with real semantic relevance.
-A.4. Source Code
-This paper is currently under review in a conference, we have submitted the source code to the venue. The source link contains the venue information so we will release it after acceptance.
+Acknowledgements.
+This work is sponsored by Tencent WeChat Rhino-Bird Focused Research Program.
+Xiangguo Sun and Hong Cheng are supported by project #MMT-p2-23 of the Shun Hing Institute of Advanced Engineering, The Chinese University of Hong Kong and by grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (No.
+CUHK 14217622).
+Chengzhi Piao is supported by grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (No.
+12202324).